OpenStudy (anonymous):

6 - 4 - 3(-2) + 6 = 12 or 2 ?

OpenStudy (anonymous):

6-4-3-2+6 =12-9 =3

OpenStudy (anonymous):

this is really easy, just do PEMDAS 6-4-*-2+6 6-4+6+6 6-4+12 2+12 it's 14